# Polynomial-Time Near-Optimal Estimation over Certain Type-2 Convex Bodies

### v1

@@ 5 Near-optimal Regression Estimation Algorithm
Since the justification of Algorithm 5 is very similar to the proof of Theorem 2.11, we omit it. However we will highlight a crucial modification that is required. We note that by a result of prasadan2025characterizingminimaxratenonparametric for ellipsoidal constraints in linear models (see Section 5.1.2 therein) we have that the minimax rate on the squared level ($\inf_{\hat{\beta}}\sup_{\beta\in B_{2}(0,\tilde{d}_{k}/2)}\mathbb{E}\|\hat{\beta}-\beta\|_{2}^{2}$) for a ball $B_{2}(0,\tilde{d}_{k}/2)$ is of the order of $n/N\wedge\tilde{d}^{2}_{k}/4$. Now, since $r<\sqrt{n}/\sqrt{N}$ (here we mean the overwritten value of $r$) and $\tilde{d}_{k}\leq 2r$ then the minimax rate reduces to $c^{\prime}\tilde{d}_{k}^{2}$, for some $c^{\prime}>0$.
@@
Suppose that $T_{2}(K)$ and $\max_{i\in[M]}\kappa(K_{(k)})$ are “small” factors and that $\log R/r$ is polynomial in $n,N$. Let $N\gtrsim n$ for an absolute constant. Then there exists a polynomial up to logarithmic factors function $\Upsilon(K)=g(T_{2}(K),\max_{i\in[M]}\kappa(K_{(k)}),\log n)$ [^20] Here we have $\log n$ in the function $g$ since the high probability in Lemma 5.2 contains an $n$ in front of the exponential which will result in a loss of logarithmic factor in $n$, i.e., we will have a term $\sqrt{(\log n)/N}$ added to the rate. We can now use the fact that $\eta^{\star}(K)\gtrsim 1/\sqrt{N}\wedge\mathrm{diam}(K)$. One can prove this similarly to the proof of Lemma 1.4 of prasadan2024some. In the case when $\mathrm{diam}(K)\lesssim 1/\sqrt{N}$ it is obvious that our estimator is minimax optimal. such that the output $\hat{\beta}$ of Algorithm 2 satisfies
@@
Since ellipsoids are type-2 sets with small type-2 constant, we note that our method provides efficient implementation for linear regression problem under ellipsoidal constraints considered in the recent paper pathak2024noisy, provided that our further assumptions on the constraint set, covariates and noise vector are satisfied.
@@ 6 Some Remarks
In this section, we give an example of a type of convex bodies that obey all conditions previously required. It has long been known donoho1990minimax that for orthosymmetric Quadratically Convex Orthosymmetric (QCO) sets, the minimax rate is given by an equation involving the Kolmogorov widths. In fact, the truncated series estimator is rate optimal over any QCO set. However, it was not previously known how one can compute a minimax optimal estimator (including the truncated series one) for QCO sets in generality. Here we argue that any orthosynmetric QCO set induces a norm which is of type-2 with small type-2 constant and is exactly 2-convex.
@@ 7 Discussion
This paper proposed the first computationally efficient algorithm for constrained mean estimation in a Gaussian sequence model setting, heavy-tailed robust mean estimation and linear regression under type-2 convex constraints, where the constraint set satisfies further regularity assumptions as outlined in Sections 2 and 3. There is one central open question that remains: can one develop near optimal polynomial-time algorithms for other (convex) constraints. Notably our algorithm does not work for $\ell_{p}$ balls for $1\leq p<2$, for which efficient near-minimax optimal soft and hard thresholding algorithms do exist johnstone2011gaussian. We further had to assume that the set $K$ is origin-symmetric. The hypothetical efficient algorithm will likely have to use a fundamentally different approach as we are no longer guaranteed that the Kolmogorov widths decay as in the origin-symmetric type-2 sets considered in this paper.
@@
Yet another question is as follows. In the work of neykov2022minimax when the constraint set is convex, one can show that the estimator is adaptive to the location of the true mean $\mu\in K$. However, we suspect that the estimator provided in this paper is not adaptive. It will be interesting to look into this rigorously: either establish that the estimator proposed here is adaptive, or refute it and provide an adaptive estimator.
Another avenue for future research is to try to remove the condition $d^{2}/\sigma^{2}\lesssim\exp(N/C_{0})/N$ in the robust algorithm. We believe this assumption is not necessary, but we will likely need to develop a new algorithm which converges without the need of the pilot estimator based on the work of depersin2022.
@@ 8 Acknowledgments
The author is grateful to Sourav Chatterjee for encouraging him to think deeply about this problem. He also benefited from valuable discussions with Andrea Montanari, Siva Balakrishnan, and Ilias Diakonikolas, as well as from exploratory conversations with ChatGPT. Finally, the author would like to thank Vijay Bhattiprolu for a helpful email exchange, where Vijay clarified several points of the central paper to our work bhattiprolu2021framework. Furthermore, the author thanks Daniel Dadush for clarifying a point in Theorem 2.5.9 in his PhD thesis.
@@ Appendix A Supplemental Proofs
This is a direct corollary of Theorem 2.5.9 of [dadush2012integer], see also [lee2018efficient] using the fact that the $\ell_{2}$ norm is $1$-Lipschitz.
@@
Step 2: Reduction to Gaussian averages. By [ledoux2013probability, Lemma 4.5],

### v2

@@ 5 Near-optimal Regression Estimation Algorithm
Since the justification of Algorithm 5 is very similar to the proof of Theorem 2.11, we omit it. However we will highlight a crucial modification that is required. We note that by a result of Prasadan and Neykov (2025a) for ellipsoidal constraints in linear models (see Section 5.1.2 therein) we have that the minimax rate on the squared level ($\inf_{\hat{\beta}}\sup_{\beta\in B_{2}(0,\tilde{d}_{k}/2)}\mathbb{E}\|\hat{\beta}-\beta\|_{2}^{2}$) for a ball $B_{2}(0,\tilde{d}_{k}/2)$ is of the order of $n/N\wedge\tilde{d}^{2}_{k}/4$. Now, since $r<\sqrt{n}/\sqrt{N}$ (here we mean the overwritten value of $r$) and $\tilde{d}_{k}\leq 2r$ then the minimax rate reduces to $c^{\prime}\tilde{d}_{k}^{2}$, for some $c^{\prime}>0$.
@@
Suppose that $T_{2}(K)$ and $\max_{i\in[M]}\kappa(K_{(k)})$ are “small” factors and that $\log R/r$ is polynomial in $n,N$. Let $N\gtrsim n$ for an absolute constant. Then there exists a polynomial up to logarithmic factors function $\Upsilon(K)=g(T_{2}(K),\max_{i\in[M]}\kappa(K_{(k)}),\log n)$ [^20] Here we have $\log n$ in the function $g$ since the high probability in Lemma 5.2 contains an $n$ in front of the exponential which will result in a loss of logarithmic factor in $n$, i.e., we will have a term $\sqrt{(\log n)/N}$ added to the rate. We can now use the fact that $\eta^{\star}(K)\gtrsim 1/\sqrt{N}\wedge\mathrm{diam}(K)$. One can prove this similarly to the proof of Lemma 1.4 of Prasadan and Neykov (2025b). In the case when $\mathrm{diam}(K)\lesssim 1/\sqrt{N}$ it is obvious that our estimator is minimax optimal. such that the output $\hat{\beta}$ of Algorithm 2 satisfies
@@
Since ellipsoids are type-2 sets with small type-2 constant, we note that our method provides efficient implementation for linear regression problem under ellipsoidal constraints considered in the recent paper Pathak et al. (2024), provided that our further assumptions on the constraint set, covariates and noise vector are satisfied.
@@ 6 Some Remarks
In this section, we give an example of a type of convex bodies that obey all conditions previously required. It has long been known Donoho et al. (1990) that for convex Quadratically Convex Orthosymmetric (QCO) sets, the minimax rate is given by an equation involving the Kolmogorov widths. In fact, the truncated series estimator is rate optimal over any QCO set. However, it was not previously known how one can compute a minimax optimal estimator (including the truncated series one) for QCO sets in generality. Here we argue that any orthosynmetric QCO set induces a norm which is of type-2 with small type-2 constant and is exactly 2-convex.
@@ 7 Discussion
This paper proposed the first computationally efficient algorithm for constrained mean estimation in a Gaussian sequence model setting, heavy-tailed robust mean estimation and linear regression under type-2 convex constraints, where the constraint set satisfies further regularity assumptions as outlined in Sections 2 and 3. There is one central open question that remains: can one develop near optimal polynomial-time algorithms for other (convex) constraints. Notably our algorithm does not work for $\ell_{p}$ balls for $1\leq p<2$, for which efficient near-minimax optimal soft and hard thresholding algorithms do exist Johnstone (2019). We further had to assume that the set $K$ is origin-symmetric. The hypothetical efficient algorithm will likely have to use a fundamentally different approach as we are no longer guaranteed that the Kolmogorov widths decay as in the origin-symmetric type-2 sets considered in this paper.
@@
Yet another question is as follows. In the work of Neykov (2022) when the constraint set is convex, one can show that the estimator is adaptive to the location of the true mean $\mu\in K$. However, we suspect that the estimator provided in this paper is not adaptive. It will be interesting to look into this rigorously: either establish that the estimator proposed here is adaptive, or refute it and provide an adaptive estimator.
Another avenue for future research is to try to remove the condition $d^{2}/\sigma^{2}\lesssim\exp(N/C_{0})/N$ in the robust algorithm. We believe this assumption is not necessary, but we will likely need to develop a new algorithm which converges without the need of the pilot estimator based on the work of Depersin and Lecué (2022).
@@ 8 Acknowledgments
The author is grateful to Sourav Chatterjee for encouraging him to think deeply about this problem. He also benefited from valuable discussions with Andrea Montanari, Siva Balakrishnan, and Ilias Diakonikolas, as well as from exploratory conversations with ChatGPT. Finally, the author would like to thank Vijay Bhattiprolu for a helpful email exchange, where Vijay clarified several points of the central paper to our work Bhattiprolu et al. (2021). Furthermore, the author thanks Daniel Dadush for clarifying a point in Theorem 2.5.9 in his PhD thesis.
@@ Appendix A Supplemental Proofs
This is a direct corollary of Theorem 2.5.9 of [5], see also [17] using the fact that the $\ell_{2}$ norm is $1$-Lipschitz.
@@
Step 2: Reduction to Gaussian averages. By [16, Lemma 4.5],